OpenStudy (nosnip):

The equation x^2 + 3px + p = 0 where p is a non-zero constant, has equal roots. Find the value of p -We have to use b^2 - 4ac

OpenStudy (nosnip):

So far I've done a=1 b=3p and c=p so that gives (3p)^2−4p=0 9p^2 - 4p = 0 p(9p−4)=0 P= 0 or 4/9 Is this correct?

OpenStudy (mathmate):

The answers and steps are both correct, congrats! :)

OpenStudy (nosnip):

brilliant thankyou very much!

OpenStudy (mathmate):

You're welcome! :)